Discuss how management utility maximization goal of a firm create principal-agent


problem? How can this problem be solved in an insurance company? A monopolist has a cost


function 200q + 15Q2


and faces a demand function given by P= 1200 – 10Q. Calculate total


revenue, marginal revenue, output and price that maximize profit-maximizing? What is its


maximal profit? Explain with the help of graphs?

1
Expert's answer
2021-12-10T15:10:27-0500

The main reasons for the principal-agent problem are conflicts of interests between two parties and the asymmetric information between them (agents tend to possess more information than principals). The principal-agent problem generally results in agency costs.

Total revenue is:

"TR = P\u00d7Q = 1200Q - 10Q^2."

Marginal revenue is:

"MR = TR'(Q) = 1200 - 20Q."

Output and price that maximize profit exist at MR = MC, so:

"MC = TC'(Q) = 200 + 20Q,"

1200 - 20Q = 200 + 20Q,

40Q = 1000,

Q = 25 units,

P = 1200 - 10×25 = 950.

Its maximal profit is:

"TP = TR - TC = 950\u00d725 - (200\u00d725 + 15\u00d715^2) = 15,375."